OpenStudy (lolacole12):

20+2(3+5) --------- 16- 49/7

OpenStudy (lolacole12):

help

OpenStudy (lolacole12):

@mathstudent55 @mathmale @KendrickLamar2014 @gabbyalicorn @gabbyalicorn @urkhada @gabbyalicorn @Michele_Laino @leahhhmorgannn

OpenStudy (lolacole12):

@Mehek14

OpenStudy (gabbyalicorn):

I'll help. Give me a sec.

OpenStudy (lolacole12):

ty

OpenStudy (lolacole12):

36/9=4

OpenStudy (gabbyalicorn):

You're welcome. :) 20+2(3+5) =36 & 16−49/7 =16/1−49/7 =16/1+−49/7 =16/1+−71 =16+−7/1 =9 So: 36/9 =4

OpenStudy (lolacole12):

one more @gabbyalicorn

OpenStudy (gabbyalicorn):

mk just make another post please.

OpenStudy (lolacole12):

ok